# Synthesize Privacy-Preserving High-Resolution Images via Private Textual Intermediaries

### v2

@@ C.1 Large Language Model and Diffusion Model APIs
SDXL-Turbo SDXL-base-1.0
Meta-Llama-3-8B-Instruct 26.71 25.42
qwen-plus 27.82 25.29
Table 4: FID (lower is better) on LSUN Bedroom with $\epsilon=1.0$. Tested on different LLM and diffusion APIs. Other experiment settings are the same as default.
@@
Appendix D Details of Experimental Configurations
@@
LSUN bedroom PE.run() num_samples_schedule 2000 2000
@@
ImageVotingNN() NNhistogram() lookahead_degree 0 -
@@
PEPopulation() initial_variation_api_fold 6 0
@@
ImageVotingNN() NearestNeighbors() lookahead_degree 8 -
lookahead_degree - 8
PEPopulation() initial_variation_api_fold 0 0
@@
wave-ui PE.run() num_samples_schedule 2000 2000
@@
ImageVotingNN() NNhistogram() lookahead_degree 0 -
@@
PEPopulation() initial_variation_api_fold 6 0
@@
Europeart PE.run() num_samples_schedule 2000 2000
@@
ImageVotingNN() NNhistogram() lookahead_degree 0 -
@@
PEPopulation() initial_variation_api_fold 6 0
@@
Sprite Fright PE.run() num_samples_schedule 1000 1000
@@
ImageVotingNN() NNhistogram() lookahead_degree 0 -
@@
PEPopulation() initial_variation_api_fold 6 0
@@
ImageVotingNN() NNhistogram() lookahead_degree 0 -
@@
PEPopulation() initial_variation_api_fold 6 0
@@
Table 6: Experiment Settings of SPTI and PE for FID evaluation of MM-Celeba-HQ dataset
@@
base learning rate $1\times 10^{-7}$ $1\times 10^{-7}$
@@ Appendix E Datasets
MM-Celeba-HQ 30,000 $256\times 256$ https://github.com/IIGROUP/MM-CelebA-HQ-Dataset

### v3

@@ B.2 FID comparison between Image Voting and Text Voting
B.3 Downstream Task
Despite the high quality of generated images in our method, we also consolidate their value in usage for downstream tasks. To be more specific, we test the classification accuracy on CelebA dataset using WRN-40-4 model. Here we provide figures of model performance during training.
Figure 26: Training performance (test accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=2000)
Figure 27: Training performance (train accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=2000)
Figure 28: Training performance (test accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=4000)
Figure 29: Training performance (train accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=4000)
Figure 30: Training performance (test accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=6000)
Figure 31: Training performance (train accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=6000)
Figure 32: Training performance (test accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=8000)
Figure 33: Training performance (train accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=8000)
Figure 34: Training performance (test accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=10000)
Figure 35: Training performance (train accuracy) on CelebA dataset with the same number of generated samples. ($\epsilon=10.0$, sample_size=10000)
Figure 36: Generated samples by SPTI ($\epsilon=10.0$, sample_size=10000)
Figure 37: Generated samples by PE image ($\epsilon=10.0$, sample_size=10000)
@@ C.1 Large Language Model and Diffusion Model APIs
SDXL-Turbo SDXL-base-1.0 Infinity
Meta-Llama-3-8B-Instruct 26.71 25.42 30.66
qwen-plus 26.65 24.44 31.28
Table 4: FID (lower is better) on LSUN Bedroom with $\epsilon=1.0$, tested using different LLM and diffusion APIs. All other settings are kept identical to the default configuration. The results indicate that while different API backends (e.g., SDXL-Turbo, SDXL-base-1.0, Infinity) and language models (e.g., Meta-Llama-3-8B-Instruct, Qwen-Plus) introduce minor variations in FID, their overall influence on the final image quality is limited. This demonstrates the robustness of our method to the choice of model and API implementation.
@@
Appendix D Explanation of Special Conditions
In the previous experiment on Cat dataset, we find our method to be less effective compared to baseline method (PE). So we conducted a contrast experiment to analyze the cause behind the phenomenon.
Firstly, we build a small sub-dataset from ImageNet100 dataset called bird dataset. This dataset consists of 100 images labeled goldfinch and 100 images labeled indigo bird, which corresponds to label id 70 and 2 in the ImageNet100 dataset. We found similar results to Cat dataset, so we take bird dataset as an alternate baseline.
Secondly, we proposed two possible reasons for the phenomenon that SPTI performs poorer than PE on Cat and bird dataset. One is that the composition of dataset is not diverse, while SPTI uses Aug-PE to generate DP data, which brings more diversity than PE image method, causing worse results. The other is that the SPTI performs naturally poorer than PE when the total number of dataset is small. We build another 3 datasets 5 according to these two reasons.
Here we reference our newly-built dataset using their composition in the following table. We only use images from ImageNet100 when building all the four dataset. The experiment settings follows the experiments on Cat dataset in Table 6. The experiment results in Table 5 shows that the reason why SPTI performs poorer than PE on Cat and bird dataset is because diversity. SPTI performs naturally poorer on dataset that is less diverse. This provides us with insight that text modality may bring more diversity on DP data generation than using image.
Table 5: FID (lower is better) comparison of SPTI and PE on Different Datasets
Dataset Composition Total Number of Images SPTI (ours) PE
goldfinch*100
+ 200 144.18 110.96
indigo bird*100
20 kinds of different birds,
200 111.84 168.39
each kind 10 images
goldfinch*1000
+ 2000 106.26 78.66
indigo bird*1000
20 kinds of different birds,
2000 84.25 172.93
each kind 100 images
Appendix E Details of Experimental Configurations
@@
LSUN bedroom
PE.run()
num_samples_schedule 2000 2000
@@
ImageVotingNN()
NNhistogram()
lookahead_degree 0 -
@@
PEPopulation()
initial_variation_api_fold 6 0
@@
ImageVotingNN()
NearestNeighbors()
lookahead_degree 4 -
lookahead_degree - 4
PEPopulation()
initial_variation_api_fold 0 0
@@
Wave-ui-25k
PE.run()
num_samples_schedule 2000 2000
@@
ImageVotingNN()
NNhistogram()
lookahead_degree 0 -
@@
PEPopulation()
initial_variation_api_fold 6 0
@@
Europeart
PE.run()
num_samples_schedule 2000 2000
@@
ImageVotingNN()
NNhistogram()
lookahead_degree 0 -
@@
PEPopulation()
initial_variation_api_fold 6 0
@@
Sprite Fright
PE.run()
num_samples_schedule 1000 1000
@@
ImageVotingNN()
NNhistogram()
lookahead_degree 0 -
@@
PEPopulation()
initial_variation_api_fold 6 0
@@
ImageVotingNN()
NNhistogram()
lookahead_degree 0 -
@@
PEPopulation()
initial_variation_api_fold 6 0
@@
Table 7: Experiment Settings of SPTI and PE for FID evaluation of $M$-Celeba-HQ dataset
@@
base learning rate $\times 10^{-7} 1×1$$\times 10^{-7}$
@@ Appendix F Datasets
$M$-Celeba-HQ 30,000 $256\times 256$ https://github.com/IIGROUP/$M$-CelebA-HQ-Dataset
CelebA 202,599 $178\times 178$ https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html